a) Determine the magnitude of the electric field in any point of an infinite plate charged with a uniform area charge density of 2 o=8C/m^2

b) A positive point charge q=+3uC is surrounded by a sphere with radius = 0.20 m centered on the charge. Find the electric flux through the sphere.


1
Expert's answer
2021-10-08T09:23:16-0400

a)


"E=\frac{8}{2\cdot8.85\cdot10^{-12}}=4.5\cdot10^{11}\frac{V}{m}"

b)


"\Phi=\frac{3\cdot10^{-6}}{8.85\cdot10^{-12}}=3.4\cdot10^{5}\ Vm"